I need help writing the equation for this problem. Please tell me why you set the problem up that way.
A plane traveled a distance of 1800 miles in three hours with the wind. The return trip takes 4 hours against the wind. Find the speed of the plane in still air and the speed of the wind.
Thanks.

Answer by checkley71(8403)

1,800/3=600 SPEED WITH THE WIND
(1,800/4=450 SPEED AGAINST THE WIND.
(600-450)/2
150/2=75 MPH IS THE SPEED OF THE WIND.
600-75=525 IS THE SPEED OF THE PLANE.
450+75=525 DITTO.